Give the equation of the line, in slope-intercept form, that passes through the points (-4,7) and (8,16)
Answer by mananth(16946)

x1 y1 x2 y2
-4 7 8 16

slope m = (y2-y1)/(x2-x1)
( 16 - 7 )/( 8 - -4 )
( 9 / 12 )
m= 3/ 4

Plug value of the slope and point ( -4 , 7 ) in
Y = m x + b
7.00 = -3 + b
b= 7 - 3
b= 4
So the equation will be
Y = 3/4 x + 4
m.ananth@hotmail.ca